# Diffusion-Based Image-to-Brain Signal Generation with Cross-Attention Mechanisms for Visual Prostheses

### v2

@@ Abstract
Visual prostheses have shown great potential in restoring vision for blind individuals. However, while researchers have successfully utilized M/EEG signals to evoke visual perceptions during the brain decoding stage of visual prostheses, the complementary process—converting images to M/EEG signals in the brain encoding stage—remains largely unexplored. Thus, we present the first image-to-brain signal (M/EEG) framework based on denoising diffusion probabilistic models enhanced with cross-attention mechanisms. Our framework consists of two key architectural components: a pre-trained CLIP visual encoder that extracts rich semantic representations from input images, and a cross-attention enhanced U-Net diffusion model that learns to reconstruct biologically plausible brain signals through iterative denoising. Unlike conventional generative models that rely on simple concatenation for conditioning, our cross-attention modules enable dynamic interaction between visual features and brain signal representations, facilitating fine-grained alignment during the generation process. Furthermore, we evaluate our framework on two multimodal datasets (THINGS-EEG2 and THINGS-MEG) to demonstrate its effectiveness in generating biologically plausible brain signals. Additionally, we pioneer the visualization of M/EEG topographies across all subjects in both datasets, providing intuitive demonstrations of intra-subject and inter-subject variations in brain signals.
@@ 3.2.2 Cross-Attention Enhanced U-Net
U-Net Encoder: The encoder processes the noisy brain signal through a series of downsampling blocks:
where represents the -th downsampling block with cross-attention mechanisms, and is the number of encoder layers.
U-Net Decoder: The decoder reconstructs the noise prediction through a series of upsampling blocks:
where represents the -th upsampling block with cross-attention mechanisms.
@@ 3.3.2 Cross-Attention Mechanism
The key innovation of our approach lies in the integration of cross-attention mechanisms within the U-Net architecture. Unlike conventional generative models that use simple concatenation or addition for conditioning, our cross-attention modules (Lin et al., 2022a) enable fine-grained alignment between visual features and brain signals. We modify the standard U-Net by incorporating cross-attention blocks in both the encoder and decoder paths. These blocks enable dynamic interaction between the brain signal representations and visual features during the denoising process. For each cross-attention layer, given the intermediate brain signal representation and visual embedding , the cross-attention is computed as:
@@ 3.5 Test Stage
During testing, we start with pure Gaussian noise and iteratively denoise it using the U-Net decoder . Given an input image, we first extract visual features:
Then, for each time step , we perform the denoising step using only the U-Net decoder with the image embeddings as conditioning:
where , is the noise variance at step , and represents the noise prediction computed using the U-Net decoder conditioned on the image embeddings . The final output represents the generated brain signal corresponding to the input image.
@@ 4.2 Experiment Details
During training, we use the complete framework shown in Figure 2 with paired image-brain signal data. For testing, we perform image-to-brain signal conversion by first extracting image embeddings via the CLIP visual encoder, then using these embeddings as conditioning information to guide the U-Net decoder through iterative denoising from Gaussian noise to generate brain signals.
Our diffusion model employs a UNet2DConditionModel from the Hugging Face diffusers library (von Platen et al., 2022) with dataset-specific configurations. For EEG signals, we use a sample size of with 63 channels and 250 temporal sampling points. For MEG signals, we configure the model with a sample size of with 271 channels and 200 temporal sampling points. Both models use 1 input and output channel, 4 downsampling and upsampling blocks with channel dimensions of . The downsampling path consists of: two DownBlock2D layers followed by two CrossAttnDownBlock2D layers, while the upsampling path includes: two CrossAttnUpBlock2D layers followed by two UpBlock2D layers. Cross-attention mechanisms are specifically integrated in the deeper layers to enable fine-grained alignment between visual features and brain signal representations (Ablation study on cross-modal learning can be seen in Appendix A.1). The cross-attention dimension is set to 768 to match the CLIP embedding dimension. We use the ViT-L/14 variant of CLIP as our visual encoder to extract rich semantic representations from input images.
@@ 4.3 Performance Evaluation
Cross-subject Generalization: From Tables 2 and 5, we observe that the cross-subject MSE averages (0.231 for EEG, 0.880 for MEG) are higher than the within-subject averages (0.217 for EEG, 0.763 for MEG), confirming the performance degradation in cross-subject scenarios. Examining cross-subject PCC results in Tables 3 and 6, we observe a significant decrease in cross-subject PCC values compared to the within-subject PCC results presented in Tables 1 and 4. These performance degradations align with our findings in Section 4.4 and Appendix A.2, which demonstrate that during object recognition tasks, brain signals from different subjects exhibit substantial variations in spatial extent and amplitude magnitudes, even when response locations remain relatively consistent. These inter-subject (cross-subject) variations pose significant challenges for developing models that can generalize effectively across individuals.
@@ 4.4 Visualization Analysis
I have plotted the topography for all subjects in both the THINGS-EEG2 and THINGS-MEG datasets, as detailed in Figures 3, 4, 5, 6, 7, and 8 (See Appendix A.2 for Figures 5, 6, 7, and 8).
@@ Appendix A Appendix
A.1 Ablation Study on Cross-Modal Learning
@@
In this section, we conduct an ablation study to evaluate the effectiveness of different cross-modal learning strategies for integrating brain signal representations with visual features. We compare three approaches: simple addition (Addition), feature concatenation (Concatenation), and cross-attention mechanism (Cross-Attention). We evaluate performance using two metrics: Mean Squared Error (MSE) and Pearson Correlation Coefficient (PCC).
Tables 7 and 9 present the MSE results for the THINGS-EEG2 and THINGS-MEG datasets, respectively. Our cross-attention approach achieves the lowest average MSE on both datasets (0.217 for EEG and 0.763 for MEG), demonstrating superior performance compared to concatenation (0.228 for EEG and 0.808 for MEG) and addition (0.227 for EEG and 0.811 for MEG) methods.
Tables 8 and 10 present the results on the THINGS-EEG2 and THINGS-MEG datasets, respectively, with all values reported as Pearson Correlation Coefficient (PCC). The cross-attention mechanism achieves the highest average PCC on both datasets (0.188 for EEG and 0.122 for MEG), outperforming both concatenation (0.180 for EEG and 0.109 for MEG) and addition (0.172 for EEG and 0.118 for MEG) methods. The consistent advantage of cross-attention across both datasets and across different evaluation metrics indicates that explicitly modeling the dynamic interactions between brain signals and visual features leads to better performance in cross-modal learning tasks.
A.2 Additional Topography Visualization

### v3

# Image-to-Brain Signal Generation for Visual Prosthesis with CLIP Guided Multimodal Diffusion Models

@@ Abstract
Visual prostheses hold great promise for restoring vision in blind individuals. While researchers have successfully utilized M/EEG signals to evoke visual perceptions during the brain decoding stage of visual prostheses, the complementary process of converting images into M/EEG signals in the brain encoding stage remains largely unexplored, hindering the formation of a complete functional pipeline. In this work, we present, to our knowledge, the first image‑to‑brain signal framework that generates M/EEG from images by leveraging denoising diffusion probabilistic models enhanced with cross‑attention mechanisms. Specifically, the proposed framework comprises two key components: a pretrained CLIP visual encoder that extracts rich semantic representations from input images, and a cross‑attention enhanced U‑Net diffusion model that reconstructs brain signals through iterative denoising. Unlike conventional generative models that rely on simple concatenation for conditioning, our cross‑attention modules capture the complex interplay between visual features and brain signal representations, enabling fine‑grained alignment during generation. We evaluate the framework on two multimodal benchmark datasets and demonstrate that it generates biologically plausible brain signals. We also present visualizations of M/EEG topographies across all subjects in both datasets, providing intuitive demonstrations of intra‑subject and inter‑subject variations in brain signals.
@@ 3.3.2 Cross-Attention Mechanism
The key innovation of our approach lies in the integration of cross-attention mechanisms within the U-Net architecture. Unlike conventional generative models that use simple concatenation or addition for conditioning, our cross-attention modules (Lin et al., 2022a) enable fine-grained alignment between visual features and brain signals. We modify the standard U-Net by incorporating cross-attention blocks in both the encoder and decoder paths. These blocks capture the complex interplay (Yang et al., 2024) between the brain signal representations and visual features during the denoising process. For each cross-attention layer, given the intermediate brain signal representation and visual embedding , the cross-attention is computed as:
@@ 3.5 Test Stage
During testing, we start with pure Gaussian noise and iteratively denoise it using our noise prediction network . Given an input image, we first extract visual features:
Then, for each time step , we perform the denoising step using the noise prediction network with the image embeddings as conditioning:
where , is the noise variance at step . The final output represents the generated brain signal corresponding to the input image.
@@ 4.2 Experiment Details
During training, we use the complete framework shown in Figure 2 with paired image-brain signal data. For testing, we perform image-to-brain signal conversion by first extracting image embeddings via the CLIP visual encoder, then using these embeddings as conditioning information to guide the noise prediction network through iterative denoising from Gaussian noise to generate brain signals.
Our diffusion model employs a UNet2DConditionModel from the Hugging Face diffusers library (von Platen et al., 2022) with dataset-specific configurations. For EEG signals, we use a sample size of with 63 channels and 250 temporal sampling points. For MEG signals, we configure the model with a sample size of with 271 channels and 200 temporal sampling points. Both models use 1 input and output channel, 4 downsampling and upsampling blocks with channel dimensions of . The downsampling path consists of: two DownBlock2D layers followed by two CrossAttnDownBlock2D layers, while the upsampling path includes: two CrossAttnUpBlock2D layers followed by two UpBlock2D layers. Cross-attention mechanisms are specifically integrated in the deeper layers to enable fine-grained alignment between visual features and brain signal representations. The cross-attention dimension is set to 768 to match the CLIP embedding dimension. We use the ViT-L/14 variant of CLIP as our visual encoder to extract rich semantic representations from input images.
@@ 4.3 Performance Evaluation
Cross-subject Generalization: From Tables 2 and 5(a), we observe that the cross-subject MSE averages (0.231 for EEG, 0.880 for MEG) are higher than the within-subject averages (0.217 for EEG, 0.763 for MEG), confirming the performance degradation in cross-subject scenarios. Examining cross-subject PCC results in Tables 3 and 5(b), we observe a significant decrease in cross-subject PCC values compared to the within-subject PCC results presented in Tables 1 and 4. These performance degradations align with our findings in Section 4.4 and Appendix A.3, which demonstrate that during object recognition tasks, brain signals from different subjects exhibit substantial variations in spatial extent and amplitude magnitudes, even when response locations remain relatively consistent. These inter-subject (cross-subject) variations pose significant challenges for developing models that can generalize effectively across individuals.
@@ 4.4 Visualization Analysis
I have plotted the topography for all subjects in both the THINGS-EEG2 and THINGS-MEG datasets, as detailed in Figures 3, 4, 6, 7, 8, and 9 (See Appendix A.3 for Figures 6, 7, 8, and 9).
@@
4.5 Cross-Modal Strategy Comparison
In this section, we conduct extensive experiments to evaluate the effectiveness of different cross-modal learning strategies for integrating brain signal representations with visual features, as shown in Figure 5. We compare three approaches: simple addition (Addition), feature concatenation (Concatenation), and cross-attention mechanism (Cross-Attention). We evaluate performance using two metrics: Mean Squared Error (MSE) and Pearson Correlation Coefficient (PCC).
EEG MSE Results and MEG MSE Results present the average MSE results for the THINGS-EEG2 and THINGS-MEG datasets, respectively. Our cross-attention approach achieves the lowest average MSE on both datasets (0.217 for EEG and 0.763 for MEG), demonstrating superior performance compared to concatenation (0.228 for EEG and 0.808 for MEG) and addition (0.227 for EEG and 0.811 for MEG) methods.
EEG PCC Results and MEG PCC Results present the average results on the THINGS-EEG2 and THINGS-MEG datasets, respectively, with all values reported as Pearson Correlation Coefficient (PCC). The cross-attention mechanism achieves the highest average PCC on both datasets (0.188 for EEG and 0.122 for MEG), outperforming both concatenation (0.180 for EEG and 0.109 for MEG) and addition (0.172 for EEG and 0.118 for MEG) methods. The consistent advantage of cross-attention across both datasets and across different evaluation metrics indicates that explicitly capturing the complex interplay between brain signals and visual features leads to better performance in cross-modal learning tasks.
@@ Appendix A Appendix
A.1 Use of Large Language Models
In the process of completing this paper, we use large language models (LLMs) for polishing the writing aspects of the paper. The conception and implementation of the ideas in this paper, the design of experiments, the selection of paper content, and other innovative aspects do not involve the use of LLMs.
A.2 Detailed Cross-Modal Strategy Comparison
@@
A.3 Additional Topography Visualization